# Adapt On-the-Go: Behavior Modulation for Single-Life Robot Deployment

### v2

@@ 5 Experimental Results
In this section, we evaluate the performance of ROAM empirically and assess how effectively it can adapt on-the-fly to new situations. Concretely, we aim to answer the following questions: (1) In simulated and real-world settings, how does ROAM compare to existing methods when given diverse prior behaviors/data and deployed in novel situations? (2) How does the additional cross-entropy term in the loss function of ROAM contribute to more successful utilization of the prior behaviors? In the remainder of this section, we describe our experimental setup and present our results on both a simulated and a real-world Go1 quadrupedal robot. For qualitative video results, see our project webpage: https://anniesch.github.io/adapt-on-the-go/.
General Experimental Setup. We use the setting of locomotion deployment to evaluate ROAM, as it is a challenging setting for adaptation, where the agent may naturally face a variety of different situations and must adapt its walking behavior on-the-fly without any additional supervision or human intervention. We use a Go1 quadruped robot from Unitree and MuJoCo [todorov2012mujoco] for simulation. We implemented all methods on top of the same state-of-the art implementation of SAC from [smith2022walk] as the base learning approach, with regularization additions following DroQ [hiraoka2021dropout], and RLPD [ball2023rlpd] for methods that do online fine-tuning. We include details and hyperparameters in Appendix A.5.
Comparisons. We evaluate ROAM along with the following prior methods: (1) RLPD Fine-tuning [ball2023rlpd], where we fine-tune a single policy using all the data from the prior behaviors with RLPD; (2) RMA [kumar2021rma], which trains a base policy and adaptation module that estimates environment info; (3) High-level Classifier (HLC), which trains a classifier on the data buffers of the pre-trained behaviors and uses it to select which behavior to use at a given state, as a representative method for those that train an additional behavior selection network, similar to work by [han2023lifelike]. We additionally consider an ablation, ROAM-NoFT, which uses the values of the prior behaviors to choose among behaviors but does not fine-tune with the additional cross-entropy loss. We give RMA access to unlimited online episode rollouts in each of the prior MDPs during pre-training, while all other methods use the same set of offline data and prior behaviors that are pre-trained in the prior MDPs. All methods are evaluated in the target MDP in a single episode without human interventions, and results are averaged across 10 trials.
@@ 5.1 Selecting Relevant Behaviors in Simulation
Setup. In our simulation experiments, we evaluate in two separate settings. The first setting simulates a situation where different joints become damaged or stuck during the robot’s lifetime. It uses 9 prior behaviors: each is a different limping behavior with a different joint frozen. In the single life, the agent must walk a total distance of 10 meters, and every 100 steps, one of the 3 remaining joints not covered in the prior data is frozen, and the agent must adapt its walking behavior to continue walking. The second setting simulates a situation where the robot encounters different friction levels on its different feet due to variation in terrain. It uses 4 different prior behaviors, each of which is trained with one of the 4 feet having low friction. During the single life, every 50 steps, the friction of one or two of the feet is changed to be lower than in the prior behaviors. To collect the prior behaviors, we train each behavior for 250k steps (first setting) or 50k steps (second setting) in the corresponding MDP, and use the last 40k steps as the prior data. We report the average and median number of steps taken to complete the task across 10 seeds along with the standard error and success rate (out of 10) in Figure 3. The agent is given a maximum of 10k steps to complete the task, and if it does not complete the task within this time, it is considered an unsuccessful trial.
@@
Results. As seen in Figure 3, ROAM outperforms all other methods in all three metrics of average and median number of steps taken to complete the task as well as overall success rate. In particular, in both settings, ROAM completes the task more than 2 times faster, in terms of average number of timesteps, compared to the next best method. Both RLPD fine-tuning and RMA struggle on both evaluation settings, especially the stiffness setting, demonstrating the importance of adapting in the space of behaviors rather than the space of actions for more efficient adaptation. RLPD and RMA perform better in the friction evaluation, where a single policy can still somewhat adapt to the various situations in the single life. On the other hand, HLC and ROAM-NoFT both struggle in the friction eval suite, demonstrating the importance of the additional cross-entropy term in the loss function, encouraging greater behavior specialization in different regions of the state space. These two methods perform better when the behaviors are already more distinguished, as in the limping eval suite, but they still struggle to adapt as efficiently as ROAM.
@@
Figure 4: Single-life Behavior Selection Accuracy. The percent of steps where different methods select a relevant behavior for the current situation. ROAM is able to choose a relevant behavior significantly more often on average when adapting to test-time situations than HLC and ROAM-noFT.
@@
Figure 5: Effect of the Cross-Entropy Loss on ID and OOD values. The average values of the different behaviors in states visited by that behavior (ID) vs states visited by other behaviors (OOD), before and after fine-tuning with the additional cross-entropy term for ROAM. ROAM is able to maintain high values of the behavior in ID states, while decreasing the value of the behavior in OOD states, which leads to better distinction between relevant behaviors at a given state.
@@
In Figures 5 and  5, we provide some additional empirical analysis of ROAM. First, in the stiffness evaluation, we plot the percent of steps where different methods select a relevant behavior during test-time deployment, where a held-out joint is frozen and a relevant behavior is one where an adjacent joint on the same leg is frozen or the same joint on an adjacent leg is frozen, and we see that ROAM is able to choose the most relevant behavior on average significantly more frequently than HLC and ROAM-noFT, which often select a behavior that is not relevant to the current situation. In addition, we record the average values of the different behaviors in states visited by that behavior (ID) vs states visited by other behaviors (OOD), before and after fine-tuning with the additional cross-entropy term for ROAM. We find that ROAM is able to effectively maintain high values of the behavior in familiar states, while decreasing the value of the behavior in unfamiliar OOD states, leading to better distinction between relevant behaviors at a given state.
@@ 5.2 Adapting on-the-Go1 in the Real World
Table 1: Results on the real Go1 robot on 3 different tasks: On all 3 tasks, across 3 trials for each method, ROAM significantly outperforms both comparisons in terms of both average wall clock time (s) and number of falls or readjustments needed to complete the task in a single life. NC (no complete) indicates that the task was not able to be successfully completed at all with the given method.
Setup. On the Go1 quadruped robot, we evaluate ROAM in a setting where we have a fixed set of five prior behaviors: walking, and four different behaviors where each of the legs has a joint frozen. We pre-train a base walking behavior in 18k steps and train the other behaviors by fine-tuning the walking behavior for an additional 3k steps with one of the joints frozen, all from scratch in the real world using the system from [smith2022walk]. During single-life deployment, we evaluate on the following three tasks: (1) Heavy Luggage: the robot must walk from a starting line to a finish line, while pulling a box that is 6.2 kg (13.6 lb) attached to one of the back legs. (2) Dynamic Luggage Load: the robot must walk from a starting line to a finish line, while adapting on-the-fly to a varying amount of weight between 2.36 kg (5.2 lb) and 4.2 kg (9.2 lb). We standardize each trial by adding and removing weight at the same distance from the start position. (3) Roller Skates: we fit the robot’s front two feet into roller skates, and the robot must adapt to its behavior to slide its forward legs and push off its back legs in order to walk to the end line. We report the average wall clock time in seconds and number of falls or readjustments needed to complete the task in a single life across 3 trials for each method. The tasks are shown in Figure 6, along with the locations of the robot before its first fall or readjustment for each method for each trial, where the red dots correspond to the Walking policy, blue to HLC, and purple to ROAM.
@@ 6 Conclusion and Future Work
We introduced Robust Autonomous Modulation (ROAM), which enables agents to rapidly adapt to changing, out-of-distribution circumstances during deployment. Our contribution lies in offering a principled, efficient way for agents to leverage pre-trained behaviors when adapting on-the-fly. On simulated and complex real-world tasks with a Go1 quadruped robot, our method achieves over 2x efficiency in adapting to new situations compared to existing methods. While ROAM offers significant advances in enabling agents to adapt to out-of-distribution scenarios, one current limitation lies in the dependency on the range of pre-trained behaviors; some scenarios may simply be too far out-of-distribution compared to the available prior behaviors. For example, an agent trained primarily in walking tasks would struggle to adapt to needing to jump over an obstacle. Future work could explore integrating ROAM into a lifelong learning framework, allowing agents to continuously expand their repertoire of behaviors, thereby increasing their adaptability to more unforeseen situations. We hope ROAM may open new possibilities for more versatile and self-reliant autonomous systems.
@@ A.1 Theoretical Analysis
In this section, we theoretically analyze ROAM to show that the additional cross-entropy loss in ROAM will lead to more suitable behaviors being chosen at each timestep. In particular, ROAM rescales the value functions of the behaviors so that they are less likely to overestimate in states that are out of distribution for that behavior. Our main result, given in Theorem A.1, is that with ROAM, for some weight $\beta>0$ on the cross-entropy term, at a given state, ROAM constrains each behavior’s value to be lower than the value of behaviors for which that state appears more frequently. This theorem gives us conservative generalization by reducing value overestimation in unfamiliar states–specifically, if at least one behavior is familiar with the current state, our chosen behavior will not have worse performance than the most familiar behavior.
@@
Let $p_{i}(s)$ denote the state visitation probability for a behavior $i$ at state $s$. For any state $s$ that is out of distribution for behavior $i$ and is in distribution for another behavior $j$, i.e. $p_{i}(s)\ll p_{j}(s)$, if $10<\beta<$ 1 is chosen to be large enough, then the value of behavior $i$ learned by ROAM will be bounded above compared to value of behavior $j$, i.e., $V_{i}(s)\leq V_{j}(s)$.
This theoretical result holds for all local extrema. Under perfect optimization, ROAM will reach a local minimum at convergence and the result would hold in this case. Our analysis implicitly assumes a simplified tabular setting in the absence of approximation and sampling errors, as done in prior theoretical analysis of deep RL algorithms, e.g. some of the analysis in [kumar2020conservative]. We take the full gradient of the TD error in the proof although in practice, only the semi-gradient is used during optimization. Regardless, the optimal value function is still a minimum of $\mathcal{L}_{\text{fine-tune}}$ and should satisfy the Bellman equation in the loss function at convergence. We also assume that for all states $s$, we have $V_{\text{freq}}(s)-V_{i}(s)=\left(R_{\text{freq}}(s)+\gamma E_{s^{\prime}}V_{\text{freq}}(s^{\prime})\right)-\left(R_{i}(s)+\gamma E_{s^{\prime}}V_{i}(s^{\prime})\right)<\eta<\infty$.
Theorem A.2.
Let $p_{i}(s)$ denote the state visitation probability for a behavior $b_{i}$ at state $s$. For any state $s$ that is out of distribution for behavior $b_{i}$ and is in distribution for another behavior $b_{j}$, i.e. $p_{i}(s)\ll p_{j}(s)$, if $01>\beta>$ 0 is chosen to be large enough, then the value of behavior $i$ learned by ROAM will be bounded above compared to value of behavior $b_{j}$, i.e. $V_{i}(s)\leq V_{j}(s)$.
@@
As a result, for any states $s$ that are out of distribution for behavior $i$ and in distribution for a different behavior $j$, if we choose $\beta$ large enough, the value learned will not overestimate the value compared to behavior $j$. Thus, at each time step, if one or more behaviors are familiar with the current state, the performance of the chosen behavior will not be much worse than its value function estimate. In this manner, ROAM adjusts value estimates based on degree of familiarity, mitigating overestimation risks. The ability to adjust the $\beta$ parameter offers a flexible framework to optimize for the behavior with the highest value at a given state, which will be at least as suitable as the most familiar behavior.
To illustrate the effect of the cross-entropy loss, consider the following example of choosing between two behaviors $i$ and $j$ at state $s$, where the true values $V^{\text{true}}_{i}(s)<V^{\text{true}}_{j}(s)$. The optimal choice is to choose behavior $j$ and for sake of this example let us choose behavior $j$ if $V^{\text{ROAM}}_{i}(s)<V^{\text{ROAM}}_{j}(s)$. There are the following four cases: (1) $p_{i}(s)<p_{j}(s)$ and the initial estimated $V_{i}(s)<V_{j}(s)$. Then with any $\beta>0$, the final $V^{\text{ROAM}}_{i}(s)<V^{\text{ROAM}}_{j}(s)$; (2) $p_{i}(s)<p_{j}(s)$ and the initial estimated $V_{i}(s)>V_{j}(s)$. Then by Theorem A.2, with large enough $\beta>0$, the final $V^{\text{ROAM}}_{i}(s)<V^{\text{ROAM}}_{j}(s)$; (3) $p_{i}(s)>p_{j}(s)$ and the initial estimated $V_{i}(s)<V_{j}(s)$. Then as long as $\beta$ is chosen to be not too large, the final $V^{\text{ROAM}}_{i}(s)<V^{\text{ROAM}}_{j}(s)$. (4) $p_{i}(s)>p_{j}(s)$ and the initial estimated $V_{i}(s)>V_{j}(s)$. This is the only case where ROAM may not be adjusted to work well, but this case poses a difficult situation for any behavior selection method.
@@ A.2 Algorithm Summary
[…]
@@
A.3 Additional Empirical Analysis
# Transitions % Timesteps Chosen Avg # Steps
5k 53.2 591.3
40k 78.4 573.8
Dynamic Friction Dynamic Stiffness
0.01 7610 +- 854 17.20% 2698 +- 844 2.92%
0.1 2082 +- 382 15.63% 1331 +- 263 8.25%
0.5 772 +- 179 11.85% 628 +- 19 12.35%
0.9 1466 +- 534 9.36% 735 +- 54 13.36%
Table 3: Sensitivity of $\beta$ and frequency of behavior switching. We find that a range of $\beta$ values give strong performance for ROAM.
Dynamic Friction Dynamic Stiffness
Avg # Steps ($\downarrow$) Avg # Steps ($\downarrow$)
RLPD 8089 $\pm$ 868 3797 $\pm$ 1029
ROAM (selection only) 1268 $\pm$ 334 1565 $\pm$ 339
ROAM 1018 $\pm$ 225 1199 $\pm$ 439
Finally, in our main experiments in Section 5, we range from choosing among 4-9 different prior behaviors. We run an additional experiment combining all the behaviors from both simulated suites, giving 13 total, and in Table 5, we find that for each suite, adding additional irrelevant behaviors did not significantly hurt performance and can even improve performance (as is the case for the stiffness eval). We hope future work may explore this at a larger scale.
Dynamic Friction Dynamic Stiffness
Avg # Steps ($\downarrow$) Avg # Steps ($\downarrow$)
ROAM (original, selected skills) 1018 $\pm$ 225 1199 $\pm$ 439
ROAM (all skills) 1141 $\pm$ 115 871 $\pm$ 137
Table 5: ROAM with irrelevant behaviors. Adding irrelevant behaviors to the set of prior behaviors did not significantly hurt performance and sometimes improves it, as seen in the stiffness evaluation.
A.4 General Experiment Setup Details

### v3

@@ 5 Experimental Results
In this section, we evaluate the performance of ROAM empirically and assess how effectively it can adapt on-the-fly to new situations. Concretely, we aim to answer the following questions: (1) In simulated and real-world settings, how does ROAM compare to existing methods when given diverse prior behaviors/data and deployed in novel situations? (2) How does the additional cross-entropy term in the loss function of ROAM contribute to more successful utilization of the prior behaviors? In the remainder of this section, we describe our experimental setup and present our results on both a simulated and a real-world Go1 quadrupedal robot. For qualitative video results, see our project webpage: https://sites.google.com/view/adapt-on-the-go/home.
General Experimental Setup. We use the setting of locomotion deployment to evaluate ROAM, as it is a challenging setting for adaptation, where the agent may naturally face a variety of different situations and must adapt its walking behavior on-the-fly without any additional supervision or human intervention. We use a Go1 quadruped robot from Unitree and MuJoCo (Todorov et al., 2012) for simulation. We implemented all methods on top of the same state-of-the art implementation of SAC from (Smith et al., 2022) as the base learning approach, with regularization additions following DroQ (Hiraoka et al., 2021), and RLPD (Ball et al., 2023) for methods that do online fine-tuning.
Comparisons. We evaluate ROAM along with the following prior methods: (1) RLPD Fine-tuning (Ball et al., 2023), where we fine-tune a single policy using all the data from the prior behaviors with RLPD; (2) RMA (Kumar et al., 2021), which trains a base policy and adaptation module that estimates environment info; (3) High-level Classifier (HLC), which trains a classifier on the data buffers of the pre-trained behaviors and uses it to select which behavior to use at a given state, as a representative method for those that train an additional behavior selection network, similar to work by (Han et al., 2023). We additionally consider an ablation, ROAM-NoCE, which uses the values of the prior behaviors to choose among behaviors but does not fine-tune with the additional cross-entropy loss. We give RMA access to unlimited online episode rollouts in each of the prior MDPs during pre-training, while all other methods use the same set of offline data and prior behaviors that are pre-trained in the prior MDPs.
@@ 5.1 Selecting Relevant Behaviors in Simulation
Setup. In our simulation experiments, we evaluate in two separate settings. The first setting simulates a situation where different joints become damaged or stuck during the robot’s lifetime. It uses 9 prior behaviors: each is a different limping behavior with a different joint frozen. In the single life, the agent must walk a total distance of 10 meters, and every 100 steps, one of the 3 remaining joints not covered in the prior data is frozen, and the agent must adapt its walking behavior to continue walking. The second setting simulates a situation where the robot encounters different friction levels on its different feet due to variation in terrain. It uses 4 different prior behaviors, each of which is trained with one of the 4 feet having low friction. During the single life, every 50 steps, the friction of one or two of the feet is changed to be lower than in the prior behaviors. To collect the prior behaviors, we train each behavior for 250k steps (first setting) or 50k steps (second setting) in the corresponding MDP, and use the last 40k steps as the prior data. The agent is given a maximum of 10k steps to complete the task, and if it does not complete the task within this time, it is considered an unsuccessful trial.
@@
Results. As seen in Figure 3, ROAM outperforms all other methods in all three metrics of average and median number of steps taken to complete the task. In particular, in both settings, ROAM completes the task more than 2 times faster, in terms of average number of timesteps, compared to the next best method. Both RLPD fine-tuning and RMA struggle on both evaluation settings, especially the stiffness setting, demonstrating the importance of adapting in the space of behaviors rather than the space of actions for more efficient adaptation. RLPD and RMA perform better in the friction evaluation, where a single policy can still somewhat adapt to the various situations in the single life. On the other hand, HLC and ROAM-NoCE both struggle in the friction eval suite, demonstrating the importance of the additional cross-entropy term in the loss function, encouraging greater behavior specialization in different regions of the state space. These two methods perform better when the behaviors are already more distinguished, as in the limping eval suite, but they still struggle to adapt as efficiently as ROAM.
@@
Figure 4: Single-life Behavior Selection Accuracy. The percent of steps where different methods select a relevant behavior for the current situation. ROAM is able to choose a relevant behavior significantly more often on average when adapting to test-time situations than HLC and ROAM-noCE.
@@
Figure 5: Effect of the Cross-Entropy Loss on ID and OOD values. The average values of the different behaviors in states visited by that behavior (ID) vs states visited by other behaviors (OOD), before and after fine-tuning with the additional cross-entropy term for ROAM. ROAM is able to maintain high values of the behavior in ID states, while decreasing the value of the behavior in OOD states.
@@
In Figures 5 and  5, we provide some additional empirical analysis of ROAM. First, in the stiffness evaluation, we plot the percent of steps where different methods select a relevant behavior during test-time deployment, where a held-out joint is frozen and a relevant behavior is one where an adjacent joint on the same leg is frozen or the same joint on an adjacent leg is frozen, and we see that ROAM is able to choose the most relevant behavior on average significantly more frequently than HLC and ROAM-noCE, which often select a behavior that is not relevant to the current situation. In addition, we record the average values of the different behaviors in states visited by that behavior (ID) vs states visited by other behaviors (OOD), before and after fine-tuning with the additional cross-entropy term for ROAM. We find that ROAM is able to effectively maintain high values of the behavior in familiar states, while decreasing the value of the behavior in unfamiliar OOD states.
@@ 5.2 Adapting on-the-Go1 in the Real World
Table 1: Results on the real Go1 robot on 3 different tasks: On all 3 tasks, across 3 trials for each method, ROAM significantly outperforms both comparisons in terms of both average wall clock time (s) and number of falls or readjustments needed to complete the task in a single life. NC (no complete) indicates that the task was not able to be successfully completed with the given method.
Setup. On the Go1 quadruped robot, we evaluate ROAM in a setting where we have a fixed set of five prior behaviors: walking, and four different behaviors where each of the legs has a joint frozen. We pre-train a base walking behavior in 18k steps and train the other behaviors by fine-tuning the walking behavior for an additional 3k steps with one of the joints frozen, all from scratch in the real world using the system from (Smith et al., 2022). During single-life deployment, we evaluate on the following three tasks: (1) Heavy Luggage: the robot must walk from a starting line to a finish line, while pulling a box that is 6.2 kg (13.6 lb) attached to one of the back legs. (2) Dynamic Luggage Load: the robot must walk from a starting line to a finish line, while adapting on-the-fly to a varying amount of weight between 2.36 kg (5.2 lb) and 4.2 kg (9.2 lb). We standardize each trial by adding and removing weight at the same distance from the start position. (3) Roller Skates: we fit the robot’s front two feet into roller skates, and the robot must adapt to its behavior to slide its forward legs and push off its back legs in order to walk to the end line. We report the average wall clock time in seconds and number of falls or readjustments needed to complete the task in a single life across 3 trials for each method. The tasks are shown in Figure 6, along with the locations of the robot before its first fall or readjustment for each method for each trial.
@@
5.3 Empirical Analysis of ROAM
# Transitions % Timesteps Chosen Avg # Steps
5k 53.2 591.3
40k 78.4 573.8
Dynamic Friction Dynamic Stiffness
0.01 7610 $\pm$ 854 17.20% 2698 $\pm$ 844 2.92%
0.1 2082 $\pm$ 382 15.63% 1331 $\pm$ 263 8.25%
0.5 772 $\pm$ 179 11.85% 628 $\pm$ 19 12.35%
0.9 1466 $\pm$ 534 9.36% 735 $\pm$ 54 13.36%
Table 3: Sensitivity of $\beta$ and frequency of behavior switching. We find that a range of $\beta$ values give strong performance.
Dynamic Friction Dynamic Stiffness
Avg # Steps ($\downarrow$) Avg # Steps ($\downarrow$)
RLPD 8089 $\pm$ 868 3797 $\pm$ 1029
ROAM (selection only) 1268 $\pm$ 334 1565 $\pm$ 339
ROAM 1018 $\pm$ 225 1199 $\pm$ 439
Dynamic Friction Dynamic Stiffness
Avg # Steps ($\downarrow$) Avg # Steps ($\downarrow$)
ROAM (original, selected skills) 1018 $\pm$ 225 1199 $\pm$ 439
ROAM (all skills) 1141 $\pm$ 115 871 $\pm$ 137
Table 5: ROAM with irrelevant behaviors. Adding irrelevant behaviors to the set of prior behaviors does not significantly hurt performance and sometimes improves it, as seen in the stiffness evaluation.
Finally, in our main experiments in Section 5, we range from choosing among 4-9 different prior behaviors. We run an additional experiment combining all the behaviors from both simulated suites, giving 13 total, and in Table 5, we find that for each suite, adding additional irrelevant behaviors did not significantly hurt performance and can even improve performance (as is the case for the stiffness eval).
@@ 6 Conclusion and Future Work
We introduced Robust Autonomous Modulation (ROAM), which enables agents to rapidly adapt to changing, out-of-distribution circumstances during deployment. Our contribution lies in offering a principled, efficient way for agents to leverage pre-trained behaviors when adapting on-the-fly. On simulated and complex real-world tasks with a Go1 quadruped robot, our method achieves over 2x efficiency in adapting to new situations compared to existing methods. While ROAM offers significant advances in enabling agents to adapt to out-of-distribution scenarios, one current limitation lies in the dependency on the range of pre-trained behaviors; some scenarios may simply be too far out-of-distribution compared to the available prior behaviors. Future work could explore integrating ROAM into a lifelong learning framework, allowing agents to continuously expand their repertoire of behaviors and increasing their adaptability to more unforeseen situations.
@@ A.1 Theoretical Analysis
We present the proof of the following theorem, discussed in Section 4.
@@
Let $p_{i}(s)$ denote the state visitation probability for a behavior $b_{i}$ at state $s$. For any state $s$ that is out of distribution for behavior $b_{i}$ and is in distribution for another behavior $b_{j}$, i.e. $p_{i}(s)\ll p_{j}(s)$, if $01>\beta>$ 0 is chosen to be large enough, then the value of behavior $b_{i}$ learned by ROAM will be bounded above compared to value of behavior $b_{j}$, i.e. $V_{i}(s)\leq V_{j}(s)$.
@@
To illustrate the effect of the cross-entropy loss, consider the following example of choosing between two behaviors $i$ and $j$ at state $s$, where the true values $V^{\text{true}}_{i}(s)<V^{\text{true}}_{j}(s)$. The optimal choice is to choose behavior $b_{j}$ and for sake of this example let us choose behavior $b_{j}$ if $V^{\text{ROAM}}_{i}(s)<V^{\text{ROAM}}_{j}(s)$. There are the following four cases: (1) $p_{i}(s)<p_{j}(s)$ and the initial estimated $V_{i}(s)<V_{j}(s)$. Then with any $\beta>0$, the final $V^{\text{ROAM}}_{i}(s)<V^{\text{ROAM}}_{j}(s)$; (2) $p_{i}(s)<p_{j}(s)$ and the initial estimated $V_{i}(s)>V_{j}(s)$. Then by Theorem A.1, with large enough $\beta>0$, the final $V^{\text{ROAM}}_{i}(s)<V^{\text{ROAM}}_{j}(s)$; (3) $p_{i}(s)>p_{j}(s)$ and the initial estimated $V_{i}(s)<V_{j}(s)$. Then as long as $\beta$ is chosen to be not too large, the final $V^{\text{ROAM}}_{i}(s)<V^{\text{ROAM}}_{j}(s)$. (4) $p_{i}(s)>p_{j}(s)$ and the initial estimated $V_{i}(s)>V_{j}(s)$. This is the only case where ROAM may not be adjusted to work well, but this case poses a difficult situation for any behavior selection method.
@@ A.2 Algorithm Summary
8:     $Q_{b^{*},\text{orig}}(s,a),\pi_{b^{*}}\leftarrow\textsc{RL}(Q_{b^{*},\text{orig}}(s,a),\pi_{b^{*}},\mathcal{D}^{b^{*}}_{\text{online}})$
@@
A.3 General Experiment Setup Details